OpenStudy (anonymous):

a , b, c are 3 consecutive terms of an A.P if tan a , tan b , tan c (b= not equal pie/2) are also in A.P then? (a) tan b = 2 tan a (b) tan a = tan b = tan c

OpenStudy (anonymous):

@kartiksriramk @dumbcow @hartnn

hartnn (hartnn):

are these 2 options? or to be proved??

OpenStudy (anonymous):

options

OpenStudy (anonymous):

nop.....The answer should be (b) tan a = tan b = tan c

hartnn (hartnn):

hmmm,working on it,wait

OpenStudy (anonymous):

@mahmit2012 @mahmit2012 @mahmit2012 @mahmit2012

hartnn (hartnn):

i got it but its a very lengthy solution..... u know that if 3 numbers are in AP and also in GP then these three are equal?? because i am getting tanA,tanB and tanC in GP that is i am getting tanA*tanC=tanB^2=3

OpenStudy (dumbcow):

there is another lengthy solution use fact that : b-a = c-b --> tan(b-a) = tan(c-b) also because tan is in AP --> tan(b) = (tan a +tan c)/2 now use difference formula for tangent and substitute in for tan(b) this will prove that tan(a) = tan(c)

hartnn (hartnn):

a+c=2b----->'X' tan a +tan c =2 tan b------>'Y' A well known result : tana*tanb*tanc=tana+tanb+tanc----->'Z' from 'Y' and 'Z', tan a tan b tan c-tan b=2 tan b or tan a*tan c= 3-------->'W' now from 'X',tan (a+c)=tan 2b \[\frac{\tan a + \tan c}{1- \tan a*\tan c}=\frac{2 \tan b}{1-(\tan b)^2}\] \[\frac{2 \tan b }{1-3}=\frac{2 \tan b}{1- (\tan b)^2}\] from 'W' and 'X' which gives (tan b)^2=3 implies,tan a* tanc = (tan b)^2 implies tan a ,tan b,tan c are also in GP and in AP also hence, tana=tanb=tanc @Yahoo